The annual sales for​ Salco, Inc. were $ 4.51 million last year. The​ firm's end-of-year balance sheet was as​ follows:

Current assets

$510,000

Liabilities

$1,016,500

Net fixed assets

1,523,000

​Owners' equity

1,016,500

Total Assets

2,033,000

Total

$2,033,000

Salco's income statement for the year was as​ follows:

Sales

$4,510,000

​Less: Cost of goods sold

(3,507,000)

Gross profit

$1,003,000

​Less: Operating expenses

(496,000)

Net operating income

$507,000

​Less: Interest expense

(91,000)

Earnings before taxes

$416,000

​Less: Taxes

​(35 %35%​)

(145,600)

Net income

$270,400

A. Calculate​ Salco's total asset​ turnover, operating profit​ margin, and operating return on assets. The​ company's total asset turnover is _ times. (Round to two decimal​ places.)


B. Salco plans to renovate one of its plants and the renovation will require an added investment in plant and equipment of $ 1.05 million. The firm will maintain its present debt ratio of 50 percent when financing the new investment and expects sales to remain constant. The operating profit margin will rise to 13.5 percent. What will be the new operating return on assets ratio​ (i.e., net operating income divided by total ​assets) for Salco after the​ plant's renovation?


C. Given that the plant renovation in part ​(b​) occurs and​ Salco's interest expense rises by $ 50 comma 000 per​ year, what will be the return earned on the common​ stockholders' investment? Compare this rate of return with that earned before the renovation. Based on this​ comparison, did the renovation have a favorable effect on the profitability of the​ firm?
Answer #1

A) Asset Turnover = Total Sales /Total assets =4,510,000/2,033,000 = 2.22

operating profit​ margin, = Operating income / Sales = 507,000/4,510,000 = 11.24%

operating return on assets = Operating income / Total assets = 507,000/ 2,033,000= 24.94%

B) The new operating return on assets for Salco after the plant renovation = Operating margin * Sales / (Total assets)

= 13.5% * (4,510,000)/(2,033,000+1,050,000) = 19.75%

C) New Net operating income = 13.5% * (4,510,000) = 608850

Earnings before taxes = New net operating income - Interest - Additional expense = 608850 - 91000 -50000 = 458850

New Net income = Earning before taxes - Taxes = 458850 - 35%*458850 = $298252.5

Return on common equity post renovation = (Net income ) / Common equity

Common equity increases by 1.05 million / 2 because 50% of asset is equity finaned.

ROCE post renovation = (298252.5) / ( 1,016,500 + 1,050,000/2) = 19.35%

Pre renovation returns = 270,400/ 1,016,500 = 26.60%

Pre-renovation return is better.